# NeRF-Aug: Data Augmentation for Robotics with Neural Radiance Fields

### v1

@@ Abstract
Training a policy that can generalize to unknown objects is a long standing challenge within the field of robotics. The performance of a policy often drops significantly in situations where an object in the scene was not seen during training. To solve this problem, we present NeRF-Aug, a novel method that is capable of teaching a policy to interact with objects that are not present in the dataset. This approach differs from existing approaches by leveraging the speed and photorealism of a neural radiance field for augmentation. NeRF-Aug both creates more photorealistic data and runs 3.83 times faster than existing methods. We demonstrate the effectiveness of our method on 4 tasks with 11 novel objects that have no expert demonstration data. We achieve an average 69.1% success rate increase over existing methods. See video results at https://nerf-aug.github.io.
@@ I Introduction
Humans have an innate ability to interact with objects they have never encountered before. For instance, a person can intuitively approach an unknown object, pick it up, and interact with it. This is in stark contrast to existing robotic systems. For a robot, even the slightest differences in shape or color from the objects seen during training can prevent the robot from achieving success. This challenge of generalization to out-of-distribution samples is a fundamental issue in machine learning and robotics.
Many prior works have explored methods to develop policies for robots that generalize to different objects. A straightforward approach is to simply collect demonstrations involving the novel object. However, this method has significant drawbacks because creating expert demonstrations is time-consuming and expensive as it requires a human to consciously control the robot’s movements. Collecting such human demonstrations is unfeasible at scale as every new object potentially requires many new demonstrations. Another approach is to use image editing tools, e.g., the latest diffusion-based image editing [1, 2, 3, 4, 5]. While these models can effectively edit images to insert new objects, they are often slow and struggle to render the exact object that will be encountered by the robot. This inaccuracy means the current object remains out of the domain of the training set which often causes these models to fail. Alternatively, some pipelines use depth images for object manipulation [6, 7]. Unfortunately, depth images in the real world suffer from noise and incompleteness [7]. This issue is exacerbated when using mounted gripper cameras, which amplify noise as they get closer to the object. Moreover, even though depth images disregard texture and color, small geometric differences between the original (training) and novel objects can still result in confusion and failure to complete the intended task.
@@
In this work, we propose NeRF-Aug, a lightweight framework that streamlines and automates data collection for a wide range of novel objects. Our goal is to generate data samples for different tasks using novel objects without collecting more human demonstrations. To enable this, we follow the image editing paradigm, but instead of relying on slow generation frameworks, we propose grounding the editing process in the scene using the 3D model of a novel object with a Neural Radiance Field (NeRF) [8] representation. We augment the training data for the robot’s policy using this edited scene. Our framework uses existing demonstrations of a different object and generates NeRF-Augmented (NeRF-Aug) synthetic data (Fig. 2) that can be used in imitation learning policies.
Figure 2: An illustration of our pipeline from beginning to end. We first train an object-level NeRF of a novel object (left). We then simultaneously erase the object in question with an off-the-shelf inpainter (top) and leverage NeRF to render images of a new object in the same position as the original object (bottom). We use the final synthetic dataset to train a new policy for the robot (right).
We demonstrate that the synthetic data generated by the NeRF-Aug framework is almost indistinguishable from real-world data. Moreover, we show that compared to existing diffusion-based image editing techniques, our method runs significantly faster, creates photorealistic images, and can consistently render objects at a wider degree of viewpoints. We test our method on a variety of real world tasks, and achieve a 69.1% increase in success rate while rendering synthetic data $3.83$ times faster than the baselines.
@@
  • We edit videos of existing demonstrations by removing the training object via in-painting and blending the NeRF render of a novel object into the inpainted image to generate a synthetic dataset for training the robot policy.
  • We demonstrate effective generalization on four diverse tasks using the generated synthetic data for training.
@@ II-A Imitation Leaning
Imitation learning is a common approach to creating robotic policies. Several approaches have been proposed that allow a policy to learn from existing successful trajectories. [1, 9, 10, 11] introduce various methods that take observations as input and predict actions. Of these, [1, 10] are multi-input models that input values from different types of data at the same time. [12, 13, 14, 15] go a step further and use an expert to actively correct the policy when as it executes trajectories. [16, 17, 18, 19] take initial demonstrations and learn a reward function which subsequently trains a reinforcement learning policy.
@@ II-B Neural Radiance Fields (NeRFs)
In recent years, there has been a surge in research on NeRFs in both computer graphics and robotics. [20, 21, 22, 23, 24] are all NeRF approaches that are capable of rendering high resolution images in real time. Other methods focus on enhancing the robustness of NeRF models against real-world constraints, such as a limited number of images [25], camera pose noise [26], and glare [27]. Finally, other novel-view synthesis techniques [28, 29, 30, 31] remove the traditional multi-layer perceptron used to volume render images with other trainable structures.
@@ II-C Synthetic Data in Robotics
Another direction for using synthetic data is using diffusion models in robotics. [1, 2, 3, 4, 5] use diffusion models to randomize the texture of objects and swap objects in a scene. [46, 47] use an action-conditioned diffusion model to simulate synthetic videos of a robot performing a task. [48] use a text-to-image diffusion model to synthesize a goal image of objects in a desired location for a goal conditioned reinforcement learning model to subsequently rearrange. [12] uses diffusion to model expert online corrections.
@@ III Preliminaries
Neural Radiance Fields. NeRF is a method used for rendering novel views of a scene or object. Formally, we assume a dataset of images and corresponding camera poses $D=\{(I_{k},T_{k})\}_{k=1}^{K}$ where $T_{k}\in\text{SE}(3)$ and $I_{k}\in\mathbb{R}^{w\times h\times 3}$. At test time, we are able to render a photorealistic view at an unknown pose $T$ through volumetric rendering. The result is a photorealistic image of a given scene through querying the NeRF model with a requested camera pose.
Figure 3: Our setup. We have two Realsense D435 cameras, angled parallel to the table. The robot arm is an UR5e from Universal Robotics.
@@ IV Our Approach: NeRF-Aug
An illustrative overview of generating synthetic data using our proposed NeRF-Aug approach is shown in Fig. 2. First, we capture multi-view images of a novel object and learn a corresponding NeRF model. Next, we use the robot arm’s gripper position to calculate the camera pose at each image in our trajectory, and we use these to query the NeRF model to render an object-level image at the same position and orientation as our original object in the training example. Then, we combine the NeRF rendering and the original image to create a new synthetic image of the robot handling a different object. Finally, we train our policy on this new dataset and evalaute it on a task consisting of the new object.
@@ IV-A Creating a NeRF model of the Novel Object
Given a novel object for which there exists no training data, we train a NeRF model for this object. Using a gripper mounted camera, we are able to collect a dataset of image-pose pairs $\left\{(I_{k},T_{k})\right\}_{k=1}^{M}$ by moving the gripper to various viewpoints both close and far away from the novel object. We use these images to train a NeRF model for the object. We found that due to real-world noise, our NeRF renders came out blurry from the millimeter-level imprecision in our measurements. To account for this, we used NerfStudio’s [49] Nerfacto model, which refines the camera intrinsic and extrinsic parameters of all images to denoise measurements.
Figure 4: A top-down view of the objects used in our real-world study. We used a variety of common household items of varying shapes, sizes and colors to show several kinds of generalization.
@@ IV-B Calculating Camera Pose relative to Object
Our method relies on accurately rendering a new object in the same position as the original training object at each frame. If the position of the object is not known ahead of time, the pose of the object can be determined by using the built-in grasp detector of a robot’s gripper. The pose of the gripper at the first frame that the object is grasped is the same as the pose of the object at the start of the trajectory. We show in Section V-F that using the gripper for calculating the object pose does not significantly decrease the success rate. Once grasped, we assume that the object does not move with respect to the gripper.
Once we know the object pose, we need to find the camera position relative to the object. Using the gripper position with respect to the world given by $T_{\text{gripper}}(t)\in\text{SE}(3)$, we are able to find the camera-to-world matrix of the gripper camera. We multiply the position of the gripper by the offset of the camera from the gripper center
$T_{\text{camera-to-world}}(t)=T_{\text{gripper}}(t)\cdot T_{\text{camera-offset}}.$ (1)
We have that the relative position of the camera with respect to the object coordinate system is denoted by
$T_{\text{camera-to-object}}(t)=T_{\text{object-to-world}}^{-1}(t)\cdot T_{\text{camera-to-world}}(t).$ (2)
@@
Figure 5: Comparison of inpainting results when replacing the original object with various objects in our training set. Instruct Pix-to-Pix and GenAug create images that show a new object, but do not recreate our specific object with appropriate dimensions and texture, this generally causes these methods to fail.
Figure 6: Where the robot is initialized and the final state of the environment for each of our tasks.
@@ IV-C NeRF Renderings
For each frame in the trajectory, we produce a rendering of our novel object at the same position and orientation as the original object by plugging our camera-to-object matrix into the NeRF model, $I_{\text{NeRF}}=\text{NeRF\_Render}\left(T_{\text{camera-to-object}}(t)\right)$. Because we are only interested in the novel object rendered from the NeRF and do not need the background of the NeRF renderings, we find a mask for the object as $M_{\text{NeRF}}$.
@@ IV-D Combining NeRF Renderings and Original Image
Once we have the NeRF rendering of the new object and the original image, we need to combine these two images. Our new object is potentially smaller than the original object, so we cannot simply copy over all the object pixels from the NeRF rendering onto our original image. If we were to do so, pixels from the original object could still remain in our new synthetic image because our new object does not completely cover the original one. Thus, we choose to erase the original object using an off-the-shelf object eraser. To do this, we use a pretrained Segment-Anything model [50] to segment the original object, Cutie video object tracker [51] to track the segmentation mask through the video frames, and ProPainter [52] to erase the object from the image using the segmentation mask. This results in a frame where the original object is removed from the frame and only the background objects/scene remains. From there on, we blend the object pixels from the NeRF renderings onto the image where the object was removed. We combine $I_{\text{NeRF}}$ and $I_{\text{no-object}}$ to get a new image with the novel object at the same pose as the original object while still keeping the background of the original image. This is done by computing
@@
Figure 7: A comparison of data augmentation results for the “pick and place” task. The top shows the original trajectory. Below that we show three methods that generate an expert synthetic trajectory for two novel objects (plastic bottle on the left, remote on the right). The last row ground-truth trajectories are created by swapping the original object (stapler) with these novel objects in the real world. GenAug is effective at swiping the texture of the object, but the color stays consistent with the original object. Instruct Pix-to-Pix is better at switching the color but also heavily changes the shading. Our method creates nearly identical trajectories to those collected in the real world.
@@
Grasping Pick and Place Drop Cup Pick Up
Methods Screwdriver Hammer Pliers Bottle Remote Eraser Spatula Green Plate Book Pan White Box
Ours 100% 90% 80% 100% 70% 40% 50% 100% 90% 100% 100%
Default 0% 0% 0% 10% 0% 0% 0% 0% 0% 0% 0%
GenAug 0% 0% 0% 0% 50% 10% 0% 0% 0% 40% 60%
Instruct Pix-to-Pix 0% 40% 30% 10% 0% 0% 0% 40% 0% 0% 30%
V Experiments
We test our method on 4 real-world tasks: grasp, pick and place, pick up, and drop. In grasp, place and place, and pick up, we swap the object that is grasped by the gripper. In drop, we switch the object that we drop a cup onto.
V-A Our Setup
In our real-world experiments, we use a Universal Robotics UR5e robot arm with a Robotiq-85 gripper. The arm is equipped with two Realsense D435 cameras with one mounted to the front of the gripper and the other mounted to the back (see Fig. 3). We run all experiments on a single Nvidia RTX A4000 GPU. We train the NeRF models with 300 images of resolution $480{\times}640$ that are taken with the robot arm, parallel to the ground. We use the default Nerfacto model from Nerf-Studio for training without any hyperparameter tuning.
For our behavior cloning model, we use BAKU [10]. The model takes image inputs from both gripper cameras to predict the next action for the robot to take. We do not allow BAKU to access gripper position for any of the trials. For the drop, pick and place, and pick up tasks, we add color jitter, which is necessary due to auto-exposure and shadows.
V-B Baselines
@@
While there has been substantial work with diffusion-based models in robot data augmentation [1, 2, 3, 4, 5], to the best of our knowledge, we are only aware of one project that made their augmentation code available to the public: GenAug [2]. As such, we are only able to compare to GenAug’s diffusion approach. We also compare to Instruct Pix-to-Pix [53], where we prompt the model to edit the image by placing a novel object at the same position as the original object.
V-C Real World Tasks
For pick and place, our initial object is a stapler which is grasped, lifted into the air, and placed on a white plate about 2 feet to the left of the stapler. An attempt is deemed successful if the object is released onto and stays on the plate. Our new objects consist of a bottle of lotion, a small remote, a dry-erase eraser, and a wooden spatula. Collecting the data requires 1.6 hours.
For pick up, our initial object is a dispenser bottle that was grasped and lifted in the air. Success is determined if the object was not in contact with the table and stayed steady within the gripper. Our novel object is a white cardboard box. The data is collected in 1.2 hours.
For grasp, our initial object is a whiteboard eraser that is grasped. Success is determined by if the object is grasped and tight so that it could not move. Our novel objects include a small hammer, an electric screwdriver, and pliers. Collecting the data takes 1.4 hours.
For drop, we drop a blue cup onto a flat object. Success is determined if the cup is successfully dropped on top of the object in question. Our initial object is a box that the robot would drop this cup onto. Our novel objects include a small black frying pan, a green plate, and a blue book. We collect the necessary data in 1.1 hours.
We show examples of start and end states for these 4 tasks in Fig. 6. We show all objects in Fig. 4. For each task, we create 200 trajectories using a hard coded expert that has access to the position of objects. To add variation to trajectories, we add seeded random noise to the output of this expert. For re-rendering, we start out with the ground truth object pose instead of calculating it.

### v2

@@ Abstract
Training a policy that can generalize to unknown objects is a long standing challenge within the field of robotics. The performance of a policy often drops significantly in situations where an object in the scene was not seen during training. To solve this problem, we present NeRF-Aug, a novel method that is capable of teaching a policy to interact with objects that are not present in the dataset. This approach differs from existing approaches by leveraging the speed, photorealism, and 3D consistency of a neural radiance field for augmentation. NeRF-Aug both creates more photorealistic data and runs 63% faster than existing methods. We demonstrate the effectiveness of our method on 5 tasks with 9 novel objects that are not present in the expert demonstrations. We achieve an average performance boost of 55.6% when comparing our method to the next best method. You can see video results at https://nerf-aug.github.io.
@@ I Introduction
Humans have an innate ability to interact with objects they have never encountered before. For instance, a person can intuitively approach an unknown object, pick it up, and interact with it. This is in stark contrast to existing robotic systems. Even the slightest differences in shape or color from the objects seen during training can prevent the robot from achieving success. This challenge of generalization to out-of-distribution samples is a fundamental issue in machine learning and robotics.
Many prior works have explored methods to develop policies for robots that generalize to different objects. A straightforward approach is to simply collect demonstrations involving the novel object. However, this method has significant drawbacks because creating expert demonstrations is time-consuming and expensive as it requires a human to consciously control the robot’s movements. Therefore, collecting such human demonstrations is infeasible at scale.
Another approach is to use image editing tools, e.g., the latest diffusion-based image editing [1, 2, 3, 4, 5]. While these models can effectively edit images to insert new objects, they are often slow and struggle to render the exact object that will be encountered by the robot. This inaccuracy means the current object remains out of the domain of the training set which often causes these models to fail. Alternatively, some pipelines use depth images for object manipulation [6, 7]. Unfortunately, depth images in the real world suffer from noise and incompleteness [7]. This issue is exacerbated when using mounted gripper cameras, which amplify noise as they get closer to the object. Moreover, even though depth images disregard texture and color, small geometric differences between the original (training) and novel objects can still result in confusion.
@@
In this work, we propose NeRF-Aug, a lightweight framework that streamlines and automates data collection for a wide range of novel objects. Our goal is to generate data samples for different tasks using novel objects without collecting more human demonstrations. To enable this, we follow the image editing paradigm, but instead of relying on slow generation frameworks, we propose using a 3D model of a novel object with a Neural Radiance Field (NeRF) [8] representation. We augment the training data for the robot’s policy using this edited scene. Our framework uses existing demonstrations of a different object and generates NeRF-Augmented (NeRF-Aug) synthetic data (Fig. 3) that can be used in imitation learning policies.
Figure 2: Comparison of inpainting results when replacing the original object with various objects in our training set. Instruct Pix-to-Pix, GenAug, and RoboAgent create images that show a new object, but do not recreate our specific object with appropriate dimensions and texture, this generally causes these methods to fail. Dreambooth has trouble with inpainting complex masks and also fails to recreate exact shade of colors.
We demonstrate that the synthetic data generated by the NeRF-Aug framework is almost indistinguishable from real-world data. Moreover, we show that compared to existing diffusion-based image editing techniques, our method runs significantly faster, creates photorealistic images, and can consistently render objects at a wide degree of viewpoints.
We test our method on a variety of real world tasks, and achieve a 55.6% increase in success rate over generation based approaches while rendering the synthetic data 63% faster than the baselines.
@@
  • We edit videos of existing demonstrations by removing the training object via in-painting and blend the NeRF render of a novel object into the inpainted image to generate a synthetic dataset.
  • We demonstrate effective generalization on five diverse tasks using the generated synthetic data for training.
Figure 3: An illustration of our pipeline from beginning to end. We first train an object-level NeRF of a novel object (left). We then simultaneously erase the object in question with an inpainter (top) and leverage NeRF to render images of a new object in the same position as the original object (bottom). We use the final synthetic dataset to train a new policy for the robot (right).
@@ II-A Imitation Leaning
Imitation learning is a common approach to creating robotic policies. Several approaches have been proposed that allow a policy to learn from existing successful trajectories. [1, 9, 10, 11] introduce various methods that take observations as input and predict actions. Of these, [1, 10] are multi-input models that input values from different types of data at the same time. [12, 13, 14, 15] go a step further and use an expert to actively correct the policy when as it executes trajectories. [16, 17, 18, 19] take initial demonstrations and learn a reward function to train a reinforcement learning policy.
@@ II-B Neural Radiance Fields (NeRFs)
In recent years, there has been a surge in research on NeRFs. [20, 21, 22, 23, 24] are all NeRF approaches that are capable of rendering high resolution images in real time. Other methods focus on enhancing the robustness of NeRF models against real-world constraints, such as a limited number of images [25], camera pose noise [26], and glare [27]. Finally, other novel-view synthesis techniques [28, 29, 30, 31] remove the traditional multi-layer perceptron used to volume render images with other trainable structures.
@@ II-C Synthetic Data in Robotics
Another direction for using synthetic data is using diffusion models in robotics. [1, 2, 3, 4, 5] use diffusion models to randomize the texture of objects and swap objects in a scene. [46, 47] use a video diffusion model to create synthetic videos of a robot performing a task and then let the robot recreate these videos.  [48] use a text-to-image diffusion model to synthesize a goal image of objects in a desired location for a goal conditioned reinforcement learning model to subsequently rearrange. [12] uses diffusion to model expert online corrections. Finally [49] uses diffusion to switch the model of robot arm performing an action in a demonstration.
@@ III Preliminaries
Neural Radiance Fields. NeRF is a method used for rendering novel views of a scene or object. Formally, we assume a dataset of images and corresponding camera poses $D=\{(I_{k},T_{k})\}_{k=1}^{K}$ where $T_{k}\in\text{SE}(3)$ and $I_{k}\in\mathbb{R}^{w\times h\times 3}$. Neural radiance fields are able to render the scene at a novel pose $T$ through volumetric rendering. The result is a photorealistic image of a given scene by querying the NeRF model with a requested camera pose.
@@ IV Our Approach: NeRF-Aug
The first step in our approach is capturing images of a novel object from multiple viewpoints and training a corresponding NeRF model. Next, we use the robot arm’s gripper position to calculate the camera pose at each image in our trajectory. Using these camera poses we query the NeRF model to render an image of the novel object at the same position and orientation as our original object in the training example. Then, we combine the NeRF rendering and the original image to create a new synthetic image of the robot handling the novel object. Finally, we train our policy on this new dataset and evaluate the original task on the novel object. An illustrative overview of our proposed NeRF-Aug approach is shown in Fig. 3.
@@ IV-A Creating a NeRF model of the Novel Object
Given a novel object for which there exists no training data, we train a NeRF model for this object. Using a gripper mounted camera, we are able to collect a dataset of image-pose pairs $\left\{(I_{k},T_{k})\right\}_{k=1}^{M}$ by moving the gripper to various viewpoints both close and far away from the novel object. We use these images to train a NeRF model for the object. In our experiments, we used NerfStudio’s [50] default Nerfacto model, which refines the camera intrinsic and extrinsic parameters of all images to denoise measurements.
@@ IV-B Calculating Camera Pose Relative to Object
Our method relies on accurately rendering a new object in the same position as the original training object at each frame. For simplicity we use a set position in this project, but the original object position can easily be calculated from the gripper position at the time the object is first grasped.
Next, NeRF-Aug finds the camera position relative to the object. Using the gripper position with respect to the world given by $T_{\text{gripper}}(t)\in\text{SE}(3)$, we are able to find the camera-to-world matrix of the gripper camera. We multiply the position of the gripper by the offset of the camera from the gripper center
$T_{\text{camera-to-world}}(t)=T_{\text{gripper}}(t)\cdot T_{\text{camera-offset}}$ (1)
The relative position of the camera with respect to the object coordinate system is denoted by
$T_{\text{camera-to-object}}(t)=T_{\text{object-to-world}}^{-1}(t)\cdot T_{\text{camera-to-world}}(t)$ (2)
@@
We also add a small amount of noise at each timestep into the camera-to-object matrix before rendering in the NeRF model. This slightly shifts the object from its’ original position to simulate a larger degree of grasp positions.
Figure 4: Our setup. We have a Realsense D455 cameras, The robot arm is a UR5e from Universal Robotics. During our tests, we use a large number of background objects such as the two dish racks and the wooden chest to reflect a real world scenario.
Figure 5: A top-down view of the objects used in our real-world study. We used a variety of common household items of varying shapes, sizes and colors to show several kinds of generalization. Left are the novel objects, right are the original objects.
@@ IV-C NeRF Renderings
For each frame in the trajectory, we produce a rendering of our novel object at the same position and orientation as the original object by plugging our camera-to-object matrix into the NeRF model, $I_{\text{NeRF}}=\text{NeRF\_Render}\left(T_{\text{camera-to-object}}(t)\right)$. Because we are only interested in the novel object rendered from the NeRF and do not need the background of the NeRF renderings, we find a mask for the object denoted as $M_{\text{NeRF}}$.
@@ IV-D Combining NeRF Renderings and Original Image
Once we have the NeRF rendering of the new object and the original image, we need to combine these two images. Our new object is potentially smaller than the original object, so we cannot simply copy over all the object pixels from the NeRF rendering onto our original image. If we were to do so, pixels from the original object could still remain in the new synthetic image. Thus, we choose to erase the original object using an inpainter. While any inpainter would work, we chose to inpaint using a NeRF model trained on the background of the scene. To do this, we use a pretrained Segment-Anything model [51] to segment the original object, Cutie video object tracker [52] to track the segmentation mask through the video frames, and a background NeRF model to switch the pixels in this mask with pixels from the background.
This results in a frame where the original object is removed from the frame and only the background objects/scene remain. From there, we blend the object pixels from the NeRF renderings onto the image where the object was removed. We combine $I_{\text{NeRF}}$ and $I_{\text{no-object}}$ to get a new image with the novel object at the same pose as the original object while still keeping the background of the original image. This is done by computing
@@
Figure 6: A comparison of data augmentation results for the “Remove Object” task. The top shows various techniques and how they compare to a original trajectory. Below that we show four methods that generate an expert synthetic trajectory for two novel objects (red spoon on the left, blue tongs on the right). GenAug is effective at swiping the texture of the object, but the color stays consistent with the original object. Instruct Pix-to-Pix is better at switching the color but also heavily changes the shading. RoboAgent adds a bunch of distractor objects that could confuse a behavior cloning model. Dreambooth inpaints an object at the same dimensions at the original object and in a different shade. Our method creates nearly identical trajectories to those collected in the real world.
V Experiments
We test our method on 5 real-world tasks: 1) Remove Object, 2) Place in Chest, 3) Serve On Plate, 4) Cook with Pot, and 5) Lift up High.
In our real-world experiments, we use a Universal Robotics UR5e robot arm with a Robotiq-85 gripper. The arm is equipped with a Realsense D455 mounted to the front of the gripper (see Fig. 4). We run all experiments on a single Nvidia RTX A4000 GPU. We use the default Nerfacto model from Nerfstudio [50] for training without any hyperparameter tuning aside from stopping the training at 4,000 timesteps for object NeRFs and 20,000 timesteps for background NeRFs.
For our behavior cloning model, we use BAKU [10]. The model takes image inputs to predict the next action for the robot to take. We do not allow BAKU to access gripper position for any of the trials. For some tasks, we also added color jitter to combat auto-exposure and shadows.
@@ V Experiments
Remove Object Place in Chest Serve on Plate Cook with Pot Lift up High
Methods Blue Tongs Red Spoon Wood Stick Swiss Miss Box White Book Large Pan Sugar Box Red Mug Hammer
Ours 70% 100% 80% 90% 60% 100% 80% 90% 100%
Default 0% 0% 0% 10% 10% 0% 0% 0% 90%
GenAug 10% 0% 0% 50% 0% 0% 0% 0% 0%
Instruct Pix-to-Pix 40% 50% 10% 50% 0% 100% 0% 0% 0%
Roboagent 30% 20% 0% 80% 60% 0% 0% 0% 10%
Dreambooth 60% 0% 60% 20% 40% 0% 0% 90% 0%
V-A Baselines
@@
While there has been substantial work with diffusion-based models in robot data augmentation [1, 2, 3, 4, 5], to the best of our knowledge, we are only aware of one project that made their augmentation code available to the public: GenAug  [2] which uses a diffusion inpainter to randomize textures and change objects. As such, we compare directly to GenAug’s diffusion approach. We also compare to a common image-editing model Instruct Pix-to-Pix [53]. In this baseline we prompt the model to edit the image by placing a novel object at the same position as the original object. We also wrote our own implementation of Roboagent [1]’s data augmentation method and compare to that. Finally, inspired by [49], we finetune a Stable Diffusion Inpainter using Dreambooth [54] to each of our novel objects with the same images used to train our NeRF models.
V-B Real World Tasks
For the Remove Object task, our initial object is a yellow spatula that rests on top of a small frying pan. The robot grasps the spatula and lifts it from the pan then drops it to the left of the pan. For this task we test on 3 novel objects: a pair of blue tongs, a red spoon, and a long wooden stick. These objects show how well NeRF-Aug can generalize to different colors and lengths of objects. For this task, we used 30 demonstrations.
For the Place in Chest task, our initial object is a mustard bottle which the robot grasps and places into a wooden chest. We tested on two novel objects: a swiss miss cardboard box and a paperback book. These novel objects differ in geometry and texture from the original object. For this task, we used 30 demonstrations.
For the Serve on Plate task, our initial object is a small white plate that we place a pear onto. We then replace the white plate with a black pan about 2 times larger than the plate. This tests how well augmentation techniques can adapt to different sizes of objects. We used 23 demonstrations.
For the Cook with Pot task, the robot removes the lid from a small pot and places it down on the table. It then picks up a stapler and places it in the pot. Our novel object is a sugar box that replaces the stapler. The sugar box has a more complicated texture than the stapler allowing us to evaluate how different textures affect policy performance as well as how well NeRF-Aug works in multi-step tasks. We use 10 demonstrations.
For the Lift up High task, the robot grasps a banana and lifts it into the air. Our novel objects were a red mug and a small yellow-black hammer in place of the banana. The mug is much wider but not as long as the banana. The hammer is much thinner but heavier than the banana. In this case, we tested how different geometries and weights can affect policy performance. We use 10 demonstrations.